OpenStudy (konradzuse):

Integration by parts. x^4 ln(x) dx u = ln(x) du = 1/x dv = x^4 v = (x^5)/5

OpenStudy (konradzuse):

\[\frac{x^5}{5} * \ln(x) - \int\limits \frac{x^5}{5} * \frac{1}{x}\]

OpenStudy (jamesj):

looking good so far!

OpenStudy (konradzuse):

\[\frac{x^5}{5} * \ln(x) - \int\limits \frac{x^4}{5}\]

OpenStudy (jamesj):

now integrate one more time

OpenStudy (konradzuse):

\[\frac{x^5}{5} * \ln(x) - \frac{x^5}{20} + c\]

OpenStudy (lgbasallote):

20?

OpenStudy (konradzuse):

I could get rid of the first x^5 with the last right? 25 I meant.

OpenStudy (jamesj):

Yes x^5/25

OpenStudy (lgbasallote):

better haha

OpenStudy (lgbasallote):

you can combine as one fraction

OpenStudy (konradzuse):

now could I do \[\ln(x) * \frac{5x^5}{25} - \frac{x^5}{25}\]

OpenStudy (lgbasallote):

dont forget the constant lol..

OpenStudy (konradzuse):

\[\ln(x) * \frac{4x^5}{25}\] + c ??

OpenStudy (konradzuse):

err tyhat messed up.

OpenStudy (lgbasallote):

hmm..it doesnt sound right...you can factor out x^5 though

OpenStudy (lgbasallote):

remember you have ln x 5 x^5 - x^5 <---you cant combine because of ln x...it's like a variable

OpenStudy (lgbasallote):

\[\large\frac{5x^5 (\ln x) - x^5 + c}{25}\] see that?

OpenStudy (konradzuse):

oh yeah I forgot I have to put the /25 under both huh...

OpenStudy (lgbasallote):

no...my point is 5x^5 (ln x) lol

OpenStudy (konradzuse):

Well is it the entire thing - x^5/25?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (konradzuse):

\[(\ln(x) * \frac{x^5}{5}) -\frac{x^5}{25} +c?\]

OpenStudy (lgbasallote):

didnt you make it 5x^5/25 to make it common denominator?

OpenStudy (konradzuse):

That's what I am being confused about.

OpenStudy (lgbasallote):

what?

OpenStudy (konradzuse):

if I made it a common denom of 25, then it would be ln(x)/5

OpenStudy (konradzuse):

I wasn't sure if the entire right side (u*v) had parenthesis or if I could deal with each part individually.

OpenStudy (lgbasallote):

hmmm let me put it this way \[\large \frac{ab}{5} = \frac{5ab}{25}\] does that help?

OpenStudy (konradzuse):

Yes I understand that.

OpenStudy (konradzuse):

In this case there isnt' an '=' though.

OpenStudy (konradzuse):

But I guess what you were saying was.

OpenStudy (konradzuse):

\[\frac{\ln(x)x^5}{5} -\frac{x^5}{25} +c? would be ) * \frac{5(\ln(x)x^5)}{25} -\frac{x^5}{25} +c?\]

OpenStudy (lgbasallote):

yes!!!

OpenStudy (konradzuse):

okies... :)

OpenStudy (konradzuse):

But before you wrote it as 5x^5 * ln(x) wouldn't it be destributed over both ln(x) and x^5? so 5ln(x) * 5x^5?

OpenStudy (lgbasallote):

\[\large ab \implies ba\] commutative property of multiplication \[\large (\ln x) x^5 \implies x^5 (\ln x)\]

OpenStudy (lgbasallote):

\[\LARGE 5(\ln x) x^5 \implies 5x^5 (\ln x)\]

OpenStudy (paxpolaris):

5ab = 5*a*b ..... NOT 5*a * 5*b

OpenStudy (lgbasallote):

it seems you are tired @KonradZuse ;) you're getting properties mixed up

OpenStudy (lgbasallote):

with all these discussion of properties...we forgot the integration question..do you know it now?